# Deep learning interpretability analysis for carbon star identification in DR3

### v1

@@ 4.2 Model interpretation and key features
The color-magnitude map shows that our new carbon star candidates occupy the location of the carbon-star golden sample but mainly concentrate in the hotter regions. From the strength map, our new carbon star candidates mainly occupy the position of weak CN band carbon stars mentioned in C2023.
The spatial distribution of the candidates is shown in Figure 17. Since the positive sample is from the golden sample of carbon stars, they are distributed throughout the Galactic inner disk, the Large Magellanic Cloud, and the Small Magellanic Cloud. However, the new candidates are distributed only in the Galactic inner disk. Therefore, the reason for their low flux values at the blue end of the spectrum is likely due to severe extinction of blue light caused by interstellar dust. This also explains why many of them became warmer ( Β‘ 2) after 3D extinction correction (Green etΒ al. 2019). Similar to the study of C-rich stars in the Milky Wayβs bar-bulge (Sanders & Matsunaga 2023), these stars are highly likely to be the result of binary evolution. However, it is also possible they are the result of single star evolution.
After cross-matching our 451 new carbon star candidates with LAMOST DR10, four common sources were obtained, two of which were labeled as carbon stars by LAMOSTβs pipeline. Cross-matching with SIMBAD yielded 73 common sources, 28 of which were labeled as C* in the main_type or other_types attribute. We will give the detailed results in Table LABEL:star_table.
@@ 5 Discussion and Analysis
According to Wallace (1962), we plotted all CN absorption line positions at the red end of the spectrum, which will be detailed in the dotted line in the following Figure 18 and Figure 20.
@@ 5.1 Cross-match with LAMOST DR10
We cross-matched the RA and Dec of all 83,028 CSTAR_XPM sample stars with the catalog of low-resolution spectra from LAMOST DR10, then we obtain the subclass types given by the LAMOSTβs pipeline (Wei etΒ al. 2014). There are 442 sources labeled as carbon stars, and 1131 are not. Among the sources not labeled as carbon stars, there are 351 G-type stars (31%), 232 K-type stars (20.5%), and 509 M-type stars (45%), the vast majority of them come from CSTAR_XPM_nonG.
@@
To better understand the formation ofcharacteristic troughs in the XP spectra of carbon stars, the LAMOST spectrum of the same source was drawn in Figure 18 to determine the absorption position of CN. Using the fast Fourier transform and interpolation method (Zhang etΒ al. 2021), the LAMOST spectrum was convolved to the resolution of the approximate XP spectra to show the difference in the position and range of the absorption region caused by different resolutions. We also plotted the same figures for common sources of G, K, and M type stars as a comparison to illustrate the areas of feature differentiation.
@@
We found that the most significant feature shared by the negative sample of G, K, and M type stars compared with the positive sample of carbon stars is the CaβII line absorption feature in the range of 8484-8662βΓ Β (Li etΒ al. 2023), which can also be identified in the XP spectra. When the temperature is low near the M type, the prominent TiO5 (7126-7135βΓ ; Li etΒ al. 2023) and KβI (7664.9-7699.0βΓ ; LΓ©pine etΒ al. 2003) absorption line features will be highlighted, which are also visible in the LAMOST and XP spectra after resolution reduction. During the change of spectral type from M to G, the absorption characteristics of TiO5 and KβI gradually weaken in the spectra until they almost disappear. This is consistent with the Figure 19 interpreted by our model. However, carbon stars exhibit broader and deeper CN absorption features at 6900-7300βΓ , and 7900-8300βΓ , and a set of three peak tops at 6830βΓ , 7780βΓ , and 9000 βΓ Β (Sanders & Matsunaga 2023). While M-type non-carbon stars also display peak structures, such as the peak top at 7050βΓ , and 7600βΓ , but their positions are entirely different. Therefore, the characteristic peaks and troughs in the spectra of carbon stars are crucial distinguishing features from other types of stars.
@@
From Figure 19, we can also see that the absorption characteristics near 8300βΓ Β and 9400βΓ Β are gradually enhanced and the sum of SHAP values increases with the temperature increase. Therefore, we hypothesize that in the CSTAR_XPM_nonG, G, and K type stars are more likely to be carbon star candidates. After a visual inspection of the 1131 stars, we estimate that G stars may contain 10% candidates and K type stars may contain 2.6% candidates. The reason is that if a low-temperature star has a high abundance of C/O, its red-end molecular band characteristics will be more obvious. It will be easier to select by the method of C2023. On the contrary, if at high-temperature conditions, the carbon molecular band characteristics would be relatively weak. In other words, if a source can still show the carbon characteristics at high temperatures, its C abundance can be considerable.
@@ 5.2 Comparison with the results of Sanders & Matsunaga (2023)
Lebzelter etΒ al. (2023) released a candidate catalog of 1,720,588 long-period variable stars (LPVs). After cross-matching with DR3 data, we found that there are 13,513 stars in the golden sample of carbon stars and only 1289 stars in CSTAR_XPM_nonG are classified as LPVs. Based on the UMAP unsupervised algorithm, Sanders & Matsunaga (2023) have presented 23,737 C-rich candidates from the LPV candidate catalog above. There are 40 of the above 1289 LPVs in the list of our 451 new carbon star candidates, and 26 are also identified as C-rich stars by Sanders & Matsunaga (2023). There are 20 common objects between our above 40 carbon candidates and 26 Carbon stars of Sanders & Matsunaga (2023), where six objects are not included in our results. We show their heat maps (see Figure 23).
@@
We can see that Gaia 963568667249950592 (the first subplot of Figure 23) is an obvious problematic spectrum, our model also assigns it a small confidence level. The spectra in subplot 2-6 of Figure 23 have relatively weak absorption features in the region around 8000βΓ Β and 9300βΓ , i.e., Feature 1 and Feature 2 in Table 4. Among these objects, Gaia 3586939158411536768 and 4056303080211083392 lack distinct CN molecular absorption band, then our model gives them lower confidence. Gaia 4314316918298388992 is weak in all Features 1-5, so the confidence level is also low. Gaia 4339500235641264256 has a stronger Feature 2, but a weaker Feature 1, resulting in a confidence level not exceeding 0.5. The molecular absorption bands of Gaia 5546572717982752640 are not prominent enough, so the confidence level is low.
@@
The majority of the golden sample of carbon stars are C-rich TP-AGB stars, typical of long-period variable stars, so it is not surprising that they are overwhelmingly classified as LPVs. Whereas the proportion of CSTAR_XPM_nonG labeled as LPV candidates is small. This may be attributed, on the one hand, they have a low signal-to-noise ratio or a lack of data points in the G time series, small variability amplitudes, or short periods. Therefore they are not detected as period signals by the SOS module of (Lebzelter etΒ al. 2023). On the other hand, they have not yet evolved to the intense third dredge-up phase of TP-AGB like LPVs, which is consistent with their location site on the HR diagram (Figure 9), or the dredge-up intensity of TP-AGB phase is insufficient.Regarding why only a few dozen of the 1289 LPVs are classified as C-rich stars, this could be due to selection effects. In CSTAR, the majority of C-rich LPVs were selected by the algorithm of C2023 and included in the golden sample, resulting in the remaining 1289 stars being predominantly O-rich LPVs. Therefore, mining possible carbon stars from CSTAR_XPM_nonG is not a simple task. Our candidates contain the vast majority of C-rich stars of Sanders & Matsunaga (2023), which is a side effect of the high completeness of our method. The other six stars are significantly different from the golden sample, but there is no denying that they still are possible carbon stars. Our model scores them with a not-low confidence level, just not above the 0.5 threshold. Although the positive sample our model learned is from the golden sample of carbon stars with strong molecular bands, it is still able to identify quite a few carbon stars exhibiting relatively weak CN molecular band features. This indicates the good generalization ability of our model and provides highly encouraging and informative results for our future systematic exploration of carbon stars in all of βs XP spectra.
@@ 5.3 Explanation of key features
The positiona of the CN absorption lines with the wavelength greater than 6900βΓ Β in the literature (Wallace 1962) were marked in the Figure 20.
From the figure, we can see that there are some differences in the position and shape of the molecular absorption band regions in the spectrum at different resolutions. From a positional perspective, the absorption regions of the Gaia carbon star spectrum are relatively more backward than that of LAMOST. From the perspective of shape, the continuous and dense CN absorption line region in the LAMOST spectrum shows the form of wide and deep CN characteristic trough in the Gaia spectrum, such as Feature 5 and Feature 2. We posit that this is the combined absorption effect of superimposed multiple molecular absorption lines in the spectrum at spectral resolution, we call it the βmixed-molecular-absorption-bandβ in the Gaia spectrum. It can be seen that Feature 1 and Feature 2 are consistent with the dense CN absorption line regions except for the small SHAP value assigned by the absorption region of Feature 5. which may be caused by the weaker signal flux values in the spectra compared to Feature 1 and Feature 2.
As for the peak observed in a spectrum, itself is a reflection of the relative decrease of opacity. When the molecular absorption on both sides of the peak becomes stronger, it will increase opacity and thus form the characteristic trough. Consequently, the region with less opacity stands out as a peak in the spectrum. Hence, the characteristic peaks of Feature 3 and Feature 4 in the shadow region are are also meaningful for distinguishing the positive and negative samples.
@@ 5.4 Reliability analysis of new candidates
From Figure 21, it is clear that there is an approximate linear relationship between the strengths and the confidence, SHAP value given by the models. Moreover, it is evident that smaller values of R773.3 and R895.0 correspond to lower confidence and SHAP values from the color-coded scatter plots. Conversely, as the values of R773.3 and R895.0 increase, so do their confidence and SHAP scores (dark color). Such strong positive correlations proved that our candidates are considered to be reliable.
The average SHAP values distribution of features is shown in Figure 22. As can be seen from this figure, the sum of average SHAP values (0.76164) of the 451 carbon star candidates is slightly lower than that of the positive sample (0.99894; see Figure 14), but much larger than that of the negative sample (0.00056), which also confirm the reliability of our candidates. We also see that the contribution of Feature 5 of the candidates almost disappears and Feature 2 becomes weaker compared to the positive sample, which is consistent with the relatively weaker carbon signatures of their spectra. While the negative feature near 8500βΓ Β (CaβII absorption lines) becomes stronger, this seems to imply that the candidates have greater metallicity (see Fig.8; Mallik 1997). It makes sense that our model treats this as a negative contribution to carbon star identification, this is because metallicity tends to be positively correlated with the oxygen levels of a starβs atmosphere (Sanders & Matsunaga 2023). Another possible reason is that the neighboring molecular bands become weaker due to temperature or C/O ratio effect, thereby highlighting the CaβII absorption line features. Although there is a slight difference in the statistical distribution of SHAP values between the new candidates and positive sample, the overall characteristics suggest that the model is more inclined to classify them as carbon stars, indicating that these candidates are more similar to the golden sample of carbon stars.
@@ 5.5 Origins of the new carbon star candidates
Investigating the origins of the 451 new carbon star candidates will be very intriguing, we try to discuss and explain. First, according to the dividing line between carbon dwarfs and carbon giants proposed by Li etΒ al. (2024), (i.e., = 5.0 mag), their luminosities are all less than 3, thus not meeting the criteria for dwarf carbon stars. Additionally, only 40 of these candidates have been reported as LPV candidates by Lebzelter etΒ al. (2023). This implies that many of them might not have evolved to the intense third dredge-up phase of the TP-AGB, and their extinction-corrected (Green etΒ al. 2019) color indicates that many of them are warmer than CN carbon stars. These reasons suggesting they are likely extrinsic carbon stars, products of binary evolution. However, it cannot be excluded that some of them are in the TP-AGB phase, but their third dredge-up event and the loss of mass are relatively mild (so warmer). Consequently, their variability amplitude is small, and they were not detected by the algorithm of Lebzelter etΒ al. (2023). The mild dredge-up intensity indirectly suggests their lower masses (Sanders & Matsunaga 2023) or that they are in the early stages of the TP-AGB phase, where the dredge-up process has not fully developed. These could explain why their spectral carbon features (such as CN and C2) are weak, although the C/O ratio has already begun to exceed 1. Of course, the higher initial oxygen abundance of the star is also a contributing factor, as can be glimpsed from the speculation on metallicity in the previous section. Finally, an accretion origin is also possible (Sanders & Matsunaga 2023), where the stars accumulate carbon by accreting surrounding carbon-rich interstellar material, thereby becoming carbon stars. Equatorial and Galactic coordinate maps indicate that they are mostly distributed in a dense, band-like region within the inner Galactic disk.
In summary, the 451 new candidates are likely external-origin carbon stars, having evolved to the giant star phase. However, the possibility of single star evolution cannot be excluded, where these carbon stars might have undergone mild third dredge-up processes or inherited carbon through accretion from surrounding interstellar material.

### v2

@@ 4.2 Model interpretation and key features
The color-magnitude map shows that our new carbon star candidates occupy the location of the carbon star golden sample but mainly concentrate in the hotter regions. From the strength map, our new carbon star candidates mainly occupy the position of weak CN band carbon stars mentioned in C2023.
The spatial distribution of the candidates is shown in Figure 17. Since the positive sample comes from the golden sample of carbon stars, they are distributed throughout the Galactic disc, the LMC, and the SMC. However, these new candidates are primarily distributed in the dense Galactic inner disc, an area characterized by substantial amounts of interstellar dust that contribute to increased extinction. Consequently, the observed degradation of their spectral shape, particularly at the blue end, is likely caused by a lower S/N or observation in crowded and highly reddened sky regions (Lebzelter etΒ al. 2023). This also explains why many of them became hotter ( Β‘ 2) after 3D extinction correction (Green etΒ al. 2019). If their color index estimates are correct, similar to the study of C-rich stars in the Milky Wayβs bar-bulge (Sanders & Matsunaga 2023), these stars are likely to be the result of binary evolution. However, the possibility of single star evolution cannot be ruled out due to the possible extinction errors.
After cross-matching our 451 new carbon star candidates with LAMOST DR10, four common sources were obtained, two of which were labeled as carbon stars by LAMOSTβs pipeline. Cross-matching with SIMBAD yielded 73 common sources, 28 of which were labeled as C* in the main_type or other_types attribute. Detailed results are provided in Table LABEL:star_table.
@@ 5 Discussion and Analysis
Following Wallace (1962), we plot all the CN absorption line positions as dotted lines at the red end of the spectrum in Figures 18 and 20.
@@ 5.1 Cross-match with LAMOST DR10
We cross-matched the RA and Dec of all the 83,028 CSTAR_XPM sample stars with the catalog of low-resolution spectra from LAMOST DR10, to obtain the subclass types given by the LAMOSTβs pipeline (Wei etΒ al. 2014). There are 442 sources labeled as carbon stars, and 1131 are not. Among the sources not labeled as carbon stars, there are 351 G-type stars (31%), 232 K-type stars (20.5%), and 509 M-type stars (45%), the vast majority of them come from CSTAR_XPM_nonG.
@@
To better understand the formation of characteristic troughs in the XP spectra of carbon stars, the LAMOST spectrum of the same source was drawn in Figure 18 to determine the CN absorption position. Using the fast Fourier transform and interpolation method of Zhang etΒ al. (2021), the LAMOST spectrum was convolved to the resolution of the approximate XP spectra to show the difference in the absorption region under different resolutions. We also plotted the same figures for common sources of G, K, and M type stars as a comparison to illustrate the areas of feature differentiation.
@@
We found that the most significant feature shared by the negative sample of G, K, and M type stars compared with the positive sample of carbon stars is the CaβII line absorption feature in the range of 8484-8662βΓ Β (Li etΒ al. 2023), which can also be identified in the XP spectra. When the temperature is low, near the M type, the prominent TiO5 (7126-7135βΓ ; Li etΒ al. 2023) and KβI (7664.9-7699.0βΓ ; LΓ©pine etΒ al. 2003) absorption line features are prominent and also visible in the LAMOST and XP spectra after resolution reduction. During the change of spectral type from M to G, the absorption characteristics of TiO5 and KβI gradually weaken in the spectra until they almost disappear. This is consistent with the Figure 19 interpreted by our model. However, carbon stars exhibit broader and deeper CN absorption features at 6900-7300βΓ , and 7900-8300βΓ , and a set of three peak tops at 6830βΓ , 7780βΓ , and 9000 βΓ Β (Sanders & Matsunaga 2023). While M-type non-carbon stars also display peak structures, such as the peak tops at 7050βΓ , and 7600βΓ , however their positions are entirely different. Therefore, both the characteristic peaks and troughs in the spectra of carbon stars are crucial distinguishing features from other types of stars.
@@
From Figure 19, we can also see that the absorption characteristics near 8300βΓ Β and 9400βΓ Β are gradually enhanced and the sum of SHAP values increases with temperature increase. Therefore, we hypothesize that in the CSTAR_XPM_nonG, G, and K type stars are more likely to be carbon star candidates. After a visual inspection of the 1131 stars, we estimate that G stars may contain 10% candidates and K type stars may contain 2.6% candidates. The reason is that if a low-temperature star has a high C abundance, its red-end molecular band characteristics will be more obvious. It will be easier to select by the method of C2023. On the contrary, if at high-temperature conditions, the carbon molecular band characteristics would be relatively weak. In other words, if a source can still show the carbon characteristics at high temperatures, its C abundance can be considerable.
@@ 5.2 Comparison with the results of Sanders & Matsunaga (2023)
Lebzelter etΒ al. (2023) released a candidate catalog of 1,720,588 long-period variables (LPVs). After cross-matching with DR3 data, we found that there are 13,513 stars in the golden sample of carbon stars and only 1289 stars in CSTAR_XPM_nonG are classified as LPVs. Based on the UMAP unsupervised algorithm, Sanders & Matsunaga (2023) have presented 23,737 C-rich candidates from the LPV candidate catalog above. There are 40 of the above 1289 LPVs in the list of our 451 new carbon star candidates, and 26 are also identified as C-rich stars by Sanders & Matsunaga (2023). There are 20 common objects between our above 40 carbon candidates and 26 Carbon stars of Sanders & Matsunaga (2023), where six objects are not included in our results. We show the heatmaps of these six object in Figure 23.
@@
We can see that Gaia 963568667249950592 (the first subplot of Figure 23) is an obvious problematic spectrum, our model also assigns it a small confidence level. The spectra in subplots 2-6 of Figure 23 have relatively weak absorption features in the region around 8000βΓ Β and 9300βΓ , i.e., Feature 1 and 2 in Table 4. Among these objects, Gaia 3586939158411536768 and 4056303080211083392 lack distinct CN molecular absorption band, then our model gives them lower confidence. Gaia 4314316918298388992 is weak in all Features 1-5, so the confidence level is also low. Gaia 4339500235641264256 has a stronger Feature 2, but a weaker Feature 1, resulting in a confidence level not exceeding 0.5. The molecular absorption bands of Gaia 5546572717982752640 are not prominent enough, so the confidence level is low.
@@
The majority of the golden sample carbon stars are TP-AGB stars, typical of long-period variable stars, so it is not surprising that they are overwhelmingly classified as LPVs by Sanders & Matsunaga (2023). However, the proportion of CSTAR_XPM_nonG stars labeled as LPV candidates is quite small. This may be attributed, to their low S/N ratio, a relative lack of data points in the time series, small variability amplitudes, or short periods. This would lead to their non-detection as period signals by the SOS module of (Lebzelter etΒ al. 2023). On the other hand, they have not yet evolved to the TP-AGB phase with periodic variability, which is consistent with their location site on the HR diagram (Figure 9). Regarding why only a few dozen of the 1289 LPVs in the CSTAR_XPM_nonG are classified as C-rich stars, this is likely due to selection effects. In CSTAR_XPM, the majority of C-rich LPVs were selected by the algorithm of C2023 and included in the golden sample, resulting in the remaining 1289 stars being predominantly O-rich LPVs, as well as some C-rich LPVs with weak carbon spectral features that were not detected by the algorithm. Therefore, mining possible carbon stars from CSTAR_XPM_nonG is not a simple task. Our candidates contain the vast majority of C-rich stars identified by Sanders & Matsunaga (2023), which demonstrates the high completeness of our method. The other six stars are significantly different from the golden sample, but there is no denying that they still are possible carbon stars. Our model scores them with a not-low confidence level, just not above the 0.5 threshold (see Section 3.1.2). Although the positive sample our model learned is from the golden sample of carbon stars with strong molecular bands, it is still able to identify quite a few carbon stars exhibiting relatively weak CN molecular band features. This demonstrates the good generalization ability of our model and encourages future systematic exploration of carbon stars across all of βs XP spectra.
@@ 5.3 Explanation of key features
The position of the CN absorption lines with a wavelength greater than 6900βΓ Β in the literature (Wallace 1962) are marked in the Figure 20.
From the figure, we can see that there are some differences in the shape of the molecular absorption band regions in the spectrum at different resolutions. The continuous and dense CN absorption line region in the LAMOST spectrum shows the form of wide and deep CN characteristic troughs in the Gaia spectrum, such as Feature 5 and Feature 2. We posit that this is the combined absorption effect of superimposed multiple molecular absorption lines in the spectrum at spectral resolution, we call it the βmixed-molecular-absorption-bandβ in the Gaia spectrum. It can be seen that Feature 1 and Feature 2 are consistent with the dense CN absorption line regions except for the small SHAP value assigned by the absorption region of Feature 5. which may be caused by the weaker signal flux values in the spectra compared to Feature 1 and Feature 2.
As for the peaks observed in the spectra, they are a reflection of the relative decrease of opacity. When the molecular absorption on both sides of the peak becomes stronger, it will increase opacity and thus form the characteristic trough. Consequently, the region with less opacity stands out as a peak in the spectrum. Hence, the characteristic peaks of Feature 3 and Feature 4 in the shadow region are are also meaningful for distinguishing the positive and negative samples.
@@ 5.4 Reliability analysis of new candidates
From Figure 21, it is clear that there is an approximately linear relationship between the strengths and the confidence, SHAP value given by the models. Moreover, it is evident that smaller values of R773.3 and R895.0 correspond to lower confidence and SHAP values from the color-coded scatter plots. Conversely, as the values of R773.3 and R895.0 increase, so do their confidence and SHAP scores (dark color). Such strong positive correlations prove that our candidates should be reliable.
The average SHAP values distribution of features is shown in Figure 22. As can be seen from this figure, the sum of average SHAP values (0.76164) of the 451 carbon star candidates is slightly lower than that of the positive sample (0.99894; see Figure 14), but much larger than that of the negative sample (0.00056), which also confirm the reliability of our candidates. We also see that the contribution of Feature 5 of the candidates almost disappears and Feature 2 becomes weaker compared to the positive sample, which is consistent with the relatively weaker carbon signatures of their spectra. While the negative feature near 8500βΓ Β (CaβII absorption lines) becomes stronger, this seems to imply that the candidates have greater metallicity (see Figure 8; Mallik 1997). It makes sense that our model treats this as a negative contribution to carbon star identification since metallicity tends to be positively correlated with the oxygen levels of a starβs atmosphere (Sanders & Matsunaga 2023). Another possible reason is that the neighboring molecular bands become weaker due to temperature or C/O ratio effect, thereby highlighting the CaβII absorption line features. Although there is a slight difference in the statistical distribution of SHAP values between the new candidates and positive sample, the overall characteristics suggest that the model is more inclined to classify them as carbon stars, indicating that these candidates are more similar to the golden sample carbon stars.
@@ 5.5 Origins of the new carbon star candidates
Since the absolute magnitudes of our new carbon star candidates are all less than 3, according to the dividing line at = 5.0 mag between carbon dwarfs and carbon giants proposed by Li etΒ al. (2024), we have selected carbon giants. Additionally, only 40 of these candidates have been reported as LPV candidates by Lebzelter etΒ al. (2023). This implies that many of them might not have evolved to the TP-AGB phase with intense TDU episodes, or alternatively, they could be TP-AGB stars but whose S/N ratio or observed amplitudes are insufficient for detection by the algorithm. The smaller amplitudes might not only be related to the evolutionary state of the stars but could also be influenced by observations in crowded regions, where the light curves of variable stars may be compressed, leading to measured amplitudes smaller than their true values (Lebzelter etΒ al. 2023), and potentially resulting in a poor S/N ratio.
Extinction correction in crowded and highly reddened sky regions must be treated with caution. If the color index corrections for the candidates are accurate, their bluer colors ( Β‘ 2) compared to the positive sample likely indicate that they are extrinsic carbon stars, i.e., products of binary evolution. Nevertheless, it cannot be excluded that some of these stars are in the TP-AGB phase. Their weaker carbon spectral features may result from temperature and metallicity effects. A higher temperature promotes the dissociation of carbon-bearing molecules in stellar atmosphere, while a higher metallicity can directly influence the C/O ratio, which determines the global spectral shape of carbon stars (Abia etΒ al. 2002). Additionally, they might be newly formed carbon stars that have undergone limited evolution (hence warmer) with few thermal pulses (TPs). Fewer TDU episodes imply that the dredged-up carbon in their atmospheres remains at relatively low levels but is already sufficient to establish a C/O ratio larger than unity (Abia etΒ al. 2002; Herwig 2005). In contrast to the evolved N-type carbon stars (C/O ratio significantly exceeding unity), that have a more pronounced mass loss (Vassiliadis & Wood 1993) and become optically obscured due to dust formation, these βyoungβ carbon stars appear to be more valuable for abundance analysis (Abia etΒ al. 2002). Furthermore, the limited amount of dredged-up carbon could also be due to the weak TDU strength in low-mass stars (Sanders & Matsunaga 2023). Finally, it is worth noting that their weak spectral carbon features may also be attributed to the lower S/N ratio or to observations in crowded or highly reddened sky regions, rather than a genuinely low carbon abundance in their atmospheres.